# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ A.3 Interpolation
Reciprocally, if we assume that $\sigma_{*}^{2}=0$, it means that there exists $x_{*}\in{\rm{argmin}}~{}f$ such that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0$. This is an expectation of nonnegative terms, such that $\nabla f_{i}(x_{*})=0$ for almost every $i\in\mathcal{I}$. Exploiting our convexity assumption, we deduce that $x_{*}$ is a minimizer of $f_{i}$ for almost every $i\in\mathcal{I}$, which shows that interpolation holds. ∎
Appendix B Reduction to a system of inequalities
B.1 Bounds from a Lyapunov decrease
@@
The first term of this Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which typically decreases for deterministic monotone gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. The standard Lyapunov for gradient descent usually contains the term $t(f(x_{t})-\min f)$, which we have replaced with the sum of the past function gaps (observe that both are of the same order in time). The last term for this Lyapunov energy is a negative cumulated sum, where the $e_{t}$’s play the role of a variance term. It is meant to compensate the fluctuations caused by the uncertainty in the SGD algorithm, and will allow the Lyapunov energy to decrease.
The core argument of our analysis is that a decrease of the Lyapunov (in expectation) yields upper bounds for SGD. Let us start by computing the discrete derivative of the energy.
Lemma B.1 (Lyapunov discrete derivative).
With our notations, for every $t=0,\dots,T-1$,
$E_{t+1}-E_{t}=a_{t+1}\|x_{t+1}-x_{*}\|^{2}-a_{t}\|x_{t}-x_{*}\|^{2}+\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}.$
Proof.
Immediate from the definition of $E_{t}$. ∎
A first result consists in deriving bounds on the function value gap, provided the Lyapunov parameter $\rho$ is nonzero. This bound will typically be obtained for convex smooth problems.
Lemma B.2 (SGD bound from Lyapunov decrease - Convex).
Consider Problem A.1. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $\rho>0$. Then
$\min\limits_{t=0,\dots,T-1}\ \mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where $\bar{e}=\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$. If we further assume that $f$ is convex, then
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
@@
$a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{0}\|^{2}}\right]\mathclose{}+\rho\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}-\sum_{t=0}^{T-1}e_{t}\sigma_{*}^{2}\leq a_{0}\|x_{0}-x_{*}\|^{2}.$
Bounding $a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{0}\|^{2}}\right]\mathclose{}\geq 0$ and dividing by $\rho T>0$ yields
$\min\limits_{t=0,\dots,T-1}\ \mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}[f(x_{t})-\min f]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\overline{e}\sigma_{*}^{2}}{\rho}.$
If $f$ is convex, we may use Jensen’s inequality to write $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]$, which implies the wanted conclusion. ∎
As a second result, a Lyapunov decrease can yield a bound of the square distance of the iterate $x_{t}$ to a minimizer, provided the Lyapunov parameter $a_{T}$ is nonzero. This is standard for strongly convex problems.
Lemma B.3 (SGD bound from Lyapunov decrease - Strongly convex).
Consider Problem A.1, and assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $a_{T}>0$. Then
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{a_{T}}+\frac{e_{T}^{sum}\sigma_{*}^{2}}{a_{T}},$
where $e_{T}^{sum}=\sum_{t=0}^{T-1}e_{t}$.
Proof.
@@
Since $f(x_{t})-\min f\geq 0$ for all $t=0,\ldots,T-1$, and $\rho\geq 0$, this yields the wanted inequality after division by $a_{T}>0$. ∎
B.2 Lyapunov decrease from a system of inequalities
Which choices of parameters $\rho,a_{0},\dots,a_{T},e_{0},\dots,e_{T-1}\geq 0$ allow the Lyapunov to decrease? The next theorem provides sufficient conditions which, as we will see later, are seemingly necessary, at least numerically. This is useful, as we will be able to focus on a system of inequalities between real numbers, instead of having to deal with all possible sequences generated by SGD. Finding parameters satisfying these sufficient conditions will be the object of the next sections.
Theorem B.4 (Sufficient conditions for Lyapunov decrease).
Let Assumptions A.4 and A.7 hold true. Assume that there exists parameters $(\alpha_{t},\beta_{t})_{t=0}^{T-1}$ such that, for every $t=0,\dots,T-1$, the following conditions are verified:
@@ B.2 Lyapunov decrease from a system of inequalities
Proof.
From Lemma B.1, we know we want to show that
$E_{t+1}-E_{t}=a_{t+1}\mathbb{E}\left[\|x_{t+1}-x_{*}\|^{2}\right]-a_{t}\mathbb{E}\left[\|x_{t}-x_{*}\|^{2}\right]+\rho\mathbb{E}\left[f(x_{t})-\min f\right]-e_{t}\sigma_{*}^{2}\leq 0.$
@@
$\|x_{t+1}-x_{*}\|^{2}=\|x_{t}-x_{*}\|^{2}+\gamma^{2}\|\nabla f_{i_{t}}(x_{t})\|^{2}-2\gamma\langle\nabla f_{i_{t}}(x_{t}),x_{t}-x_{*}\rangle.$
Therefore,
$E_{t+1}-E_{t}$ $=$ $\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}+(a_{t+1}-a_{t})\|x_{t}-x_{*}\|^{2}$
$+a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}-2a_{t+1}\gamma\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle.$
To show that this random variable is nonpositive in expectation, we shall make use of an assumption on the loss functions. While it would be natural to exploit the fact that the $f_{i}$ are $\mu$-convex and smooth for all $i\in\mathcal{I}$, our numerical simulations suggest that all we need is (SEC).
In what follows, we denote by $\mathbb{E}_{t}\coloneqq\mathbb{E}\mathopen{}\left[{\cdot\ |\ x_{0},\dots,x_{t}}\right]\mathclose{}$ the conditional expectation with respect to the iterates up to and including $x_{t}$. Property (SEC) reads
$\frac{1}{2}\mathbb{E}_{t}\mathopen{}\left[{\|\nabla f_{i}(x_{*})-\nabla f_{i}(x_{t})\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|x_{*}-x_{t}\|^{2}\leq(L-\mu)\left(\min f-f(x_{t})\right)-L\langle\nabla f(x_{t}),x_{*}-x_{t}\rangle$
at the pair $(x_{t},x_{*})$, and
$\frac{1}{2}\mathbb{E}_{t}\mathopen{}\left[{\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|x_{t}-x_{*}\|^{2}\leq(L-\mu)\left(f(x_{t})-\min f\right)+\mu\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle$
at the pair $(x_{*},x_{t})$. Multiplying these inequalities by $2\alpha_{t}\geq 0$ and $2\beta_{t}\geq 0$ respectively, and summing them gives
@@
Adding $\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$ on both sides and using the expression obtained in (B.2), we get
$\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$$\leq$ $\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$
+ $+\ 2(L-\mu)(\beta_{t}-\alpha_{t})(f(x_{t})-\min f)+2(\alpha_{t}L+\beta_{t}\mu)\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle$
$-\ (\alpha_{t}+\beta_{t})\mathbb{E}_{t}\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}-\mu L(\alpha_{t}+\beta_{t})\|x_{t}-x_{*}\|^{2}$
$=$ $\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}+(a_{t+1}-a_{t})\|x_{t}-x_{*}\|^{2}$
$+\ a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}-2a_{t+1}\gamma\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle$
+ $+\ 2(L-\mu)(\beta_{t}-\alpha_{t})(f(x_{t})-\min f)+2(\alpha_{t}L+\beta_{t}\mu)\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle$
$-\ (\alpha_{t}+\beta_{t})\mathbb{E}_{t}\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}-\mu L(\alpha_{t}+\beta_{t})\|x_{t}-x_{*}\|^{2}$
$=$ $(\rho+2(L-\mu)(\beta_{t}-\alpha_{t}))(f(x_{t})-\min f)-e_{t}\sigma_{*}^{2}$
$+(a_{t+1}-a_{t}-\mu L(\alpha_{t}+\beta_{t}))\|x_{t}-x_{*}\|^{2}+\ a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}$
$+2(\alpha_{t}L+\beta_{t}\mu-a_{t+1}\gamma)\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle$
$-(\alpha_{t}+\beta_{t})\mathbb{E}_{t}\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}.$
Taking the expectation and developing the last square yields
@@
$+(a_{t+1}-a_{t}-\mu L(\alpha_{t}+\beta_{t}))\mathbb{E}\left[\|x_{t}-x_{*}\|^{2}\right]$
$+\ (a_{t+1}\gamma^{2}-\alpha_{t}-\beta_{t})\mathbb{E}\left[\|\nabla f_{i}(x_{t})\|^{2}\right]$
$-(e_{t}+\alpha_{t}+\beta_{t})\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$
$+2(\alpha_{t}L+\beta_{t}\mu-a_{t+1}\gamma)\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle\right]$
$+2(\alpha_{t}+\beta_{t})\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),\nabla f_{i}(x_{*})\rangle\right].$
For the Lyapunov energy to be decreasing in expectation, it is enough that the right-hand side is nonpositive. For the first term, because $f(x_{t})-\min f\geq 0$, it is enough to assume that $\rho\leq 2(L-\mu)(\alpha_{t}-\beta_{t})$, which corresponds to Condition 2. For the other terms, we enforce that
$A\mathbb{E}\left[\|x_{t}-x_{*}\|^{2}\right]+B\mathbb{E}\left[\|\nabla f_{i}(x_{t})\|^{2}\right]+B^{\prime}\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ (2)
$+$ $2C\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle\right]+2D\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),\nabla f_{i}(x_{*})\rangle\right]\geq 0,$
@@
The quantity in (2) is simply the expectation of a quadratic polynomial in $X=x_{t}-x_{*}$, $Y=\nabla f_{i}(x_{t})$ and $Y^{\prime}=\nabla f_{i}(x_{*})$. Using elementary linear algebra, it can be shown that for the expression in (2) to be nonnegative, it is enough to require
@@
We postpone the proof of this claim to Lemma B.6.
A simple calculation shows that these conditions correspond exactly to the remaining wanted. In fact, $A\geq 0$ corresponds to Condition 3 and $B\geq 0$ to Condition 4. The inequality $B^{\prime}\geq 0$ is trivially satisfied because $e_{t},\alpha_{t},\beta_{t}\geq 0$ by Condition 1. Moreover, $C^{2}\leq AB$ may be developed as
$(a_{t+1}\gamma-\alpha_{t}L-\beta_{t}\mu)^{2}\leq(\mu L(\alpha_{t}+\beta_{t})a_{t}-a_{t+1})(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2}),$
which corresponds to Condition 5. Finally, $D^{2}\leq BB^{\prime}$ expands to
$(\alpha_{t}+\beta_{t})^{2}\leq(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})(e_{t}+\alpha_{t}+\beta_{t}),$
which can be simplified to
$a_{t+1}\gamma^{2}(\alpha_{t}+\beta_{t})\leq(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})e_{t},$
which is exactly Condition 6. ∎
Remark B.5 (Relaxing (SEC) when $\beta_{t}=0$).
Looking at how we used (SEC) in the above proof, it is clear that if we impose $\beta_{t}=0$, then one of the two inequalities in (SEC) is useless, and that all we need in that case is (EC). This will be, for instance, the case when proving our bounds for short step-sizes, see later in Remark C.3.
We end this section by proving the technical lemma at the core of the previous proof.
Lemma B.6 (Nonnegative quadratic polynomial).
@@
For (5) to have a solution, the above must hold simultaneously with $-\sqrt{AB^{\prime}}\leq\theta\leq\sqrt{AB^{\prime}}$. This is equivalent to the intersection of two intervals to be nonempty, which is the case if

### v2

@@ 2 Preliminaries
Reciprocally, if we assume that $\sigma_{*}^{2}=0$, it means that there exists $x_{*}\in{\rm{argmin}}~f$ such that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0$. This is an expectation of nonnegative terms, so we deduce that $\nabla f_{i}(x_{*})=0$ for almost every $i\in\mathcal{I}$. Exploiting our convexity assumption, we deduce that $x_{*}$ is a minimizer of $f_{i}$ for almost every $i\in\mathcal{I}$, which shows that interpolation holds.
3 Reducing a Lyapunov Analysis to the Resolution of a System of Inequalities
Our main approach to derive bounds for SGD is to rely on an analysis of a new parametrized family of Lyapunov energies. After introducing these Lyapunov energies, we will show how to deduce bounds for SGD from a Lyapunov decrease in Section 3.1. In a second time, we present our main technical result which states that obtaining such Lyapunov decrease can be reduced to solving a system of inequalities (see Section 3.2).
3.1 Bounds for SGD from a Lyapunov Decrease
@@
The first term of this Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which typically decreases for deterministic monotone gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. The standard Lyapunov for gradient descent usually contains the term $t(f(x_{t})-\min f)$, and such term was also considered for analyzing SGD in taylor_stochastic_2019 . Here we make a slightly different choice, replacing this term with the sum of the past function gaps (observe that both are of the same order in time). The last term for this Lyapunov energy is a negative cumulated sum, where the $e_{t}$’s measure the contribution of the variance. This term is meant to compensate the fluctuations caused by the uncertainty in the SGD algorithm, and will allow the Lyapunov energy to decrease.
A first result consists in deriving bounds on the average function value gap, provided the Lyapunov parameter $\rho$ is nonzero. This bound will typically be obtained for convex smooth problems. The proof is based on a telescopic sum and an application of Jensen’s inequality, as is often the case for convex smooth objectives, as done for instance in taylor_stochastic_2019 ; gower_sgd_2021 .
Lemma 7(Bound from Lyapunov decrease. Convex case)
Consider Problem 1, assume that $f$ is convex, and let $T\geq 1$. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $\rho>0$. Then, if we note $\bar{e}=\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$ and $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho}.$
Proof
@@
$a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}+\rho\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}-\sum_{t=0}^{T-1}e_{t}\sigma_{*}^{2}\leq a_{0}\|x_{0}-x_{*}\|^{2}.$
Bounding $a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\geq 0$ and dividing by $\rho T>0$ yields
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}[f(x_{t})-\min f]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where in the first inequality we used Jensen’s inequality.
As a second result, a Lyapunov decrease can yield a bound of the square distance of the iterate $x_{t}$ to a minimizer, provided the Lyapunov parameter $a_{T}$ is nonzero. This is standard for strongly convex problems, as we will see in Section 5. The proof, also based on a telescopic argument, is again typical for this class of problems, see for instance gower_sgd_2019 .
Lemma 8(Bound from Lyapunov decrease. Strongly convex case)
Consider Problem 1, and let $T\geq 1$. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $a_{T}>0$. Then, if we note $e_{T}^{sum}=\sum_{t=0}^{T-1}e_{t}$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{a_{T}}+\frac{e_{T}^{sum}\sigma_{*}^{2}}{a_{T}}.$
Proof
@@
Since $f(x_{t})-\min f\geq 0$ for all $t=0,\ldots,T-1$, and $\rho\geq 0$, this yields the wanted inequality after division by $a_{T}>0$.
3.2 Lyapunov Decrease from a System of Inequalities
Lemmas 7 and 8 reduce the problem of finding upper bounds for SGD to finding parameters $\rho,a_{0},\dots,a_{T},e_{0},\dots,e_{T-1}\geq 0$ that allow the Lyapunov energy to decrease. The next theorem provides sufficient conditions for this to be the case, and will be at the heart of our subsequent results. Indeed, this will allow to shift the focus from sequences generated by SGD to a system of inequalities on real numbers.
Theorem 3.1(Sufficient conditions for Lyapunov decrease)
Let Assumptions 1 and 2 hold true, and let $T\geq 1$. Assume that there exist parameters $(\alpha_{t},\beta_{t})_{t=0}^{T-1}$ such that, for every $t=0,\dots,T-1$, the following conditions are verified:
@@ 3.2 Lyapunov Decrease from a System of Inequalities
Proof
@@
$E_{t+1}-E_{t}=\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}+(a_{t+1}-a_{t})\|x_{t}-x_{*}\|^{2}+a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}-2a_{t+1}\gamma\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle.$ (1)
In what follows, we denote by $\mathbb{E}_{t}\coloneqq\mathbb{E}\mathopen{}\left[{\cdot\ |\ x_{0},\dots,x_{t}}\right]\mathclose{}$ the conditional expectation with respect to the iterates up to and including $x_{t}$. Multiplying the expected cocoercivity inequalities (Lemma 1) for the pair $(x_{t},x_{*})$ by $2\alpha_{t}\geq 0$ and for the pair $(x_{*},x_{t})$ by $2\beta_{t}\geq 0$, and summing them gives
@@
Adding $\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$ on both sides and using the expression obtained in (1), taking the expectation and developing the squares yields
@@
$+\mathbb{E}\mathopen{}\left[{A{\|x_{t}-x_{*}\|^{2}}+B{\|\nabla f_{i}(x_{t})\|^{2}}+B^{\prime}{\|\nabla f_{i}(x_{*})\|^{2}}+2C{\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle}+2D{\langle\nabla f_{i}(x_{t}),\nabla f_{i}(x_{*})\rangle,}}\right]\mathclose{}$
@@
For the Lyapunov energy to be decreasing in expectation, it is enough that the right-hand side is nonpositive. For the first term, because $f(x_{t})-\min f\geq 0$, it is enough to assume that $\rho\leq 2(L-\mu)(\alpha_{t}-\beta_{t})$, which corresponds to Condition 2. The remaining terms are simply the expectation of a quadratic polynomial in $X=x_{t}-x_{*}$, $Y=\nabla f_{i}(x_{t})$ and $Y^{\prime}=\nabla f_{i}(x_{*})$. Using elementary linear algebra, it can be shown (see Lemma 9 below) that for this polynomial to be nonnegative, it is enough to require
@@
A simple calculation shows that these conditions correspond exactly to the remaining conditions of our theorem.
Remark 1
We note that Theorem 3.1, which is the key theorem for all of our subsequent results, holds provided that the inequality of Lemma 1 holds for any pair of points $(x,x_{*})$ and $(x_{*},x)$, where $x\in\mathcal{H}$. As such, we could replace Assumption 1 by such a weaker version, which includes a class of nonconvex functions. Moreover, in the case where $\beta_{t}\equiv 0$, it is only necessary for the conclusion of Lemma 1 to hold for any pair $(x,x_{*})$, where $x\in\mathcal{H}$. We note the similarities with the recently introduced expected smoothness assumption gower_stochastic_2021 . Although this nonconvex extension is of interest, we should not delve further into it in this work.
We end this section by proving the technical lemma used the previous proof.
Lemma 9(Nonnegative quadratic polynomial)
@@
For (4) to have a solution, the above must hold simultaneously with $\sqrt{AB^{\prime}}\leq\theta\leq\sqrt{AB^{\prime}}$. This is equivalent to the intersection of two intervals to be nonempty, which is the case if